OpenStudy (anonymous):

5/3x-7/3=x solve for X.

OpenStudy (kamille):

hey, is it like this \[\frac{ 5 }{ 3x}-\frac{ 7 }{ 3 }=x\]

OpenStudy (anonymous):

5/3x-7/3=x 5/3x - x = 7/3 5/3x - 3/3x = 7/3 2/3x = 7/3 x = 7/3 times 3/2 x = 21/6

OpenStudy (anonymous):

sorry

OpenStudy (kamille):

or \[\frac{ 5 }{ \frac{ 3x-7 }{ 3 } }=x\]

OpenStudy (anonymous):

x = 7/3 times 3/2 x = 7/2

OpenStudy (anonymous):

just typo :)

OpenStudy (anonymous):

its \[\frac{ 5 }{ 3 }x\]

OpenStudy (anonymous):

Sorry I should have clarified

OpenStudy (anonymous):

i believe it would be \[\frac{ 5 }{ 3 }x - \frac{ 7}{ 3 } = x\]

OpenStudy (anonymous):

\[\frac{ 5 }{ 3 }x - \frac{ 7 }{ 3 } = X\]

OpenStudy (kamille):

oh ok, I think @gerryliyana has already solved it then

OpenStudy (anonymous):

5/3x-7/3=x 5/3x - x = 7/3 5/3x - 3/3x = 7/3 2/3x = 7/3 x = 7/3 times 3/2 x = 7/2